# Bipartite Randomized Response Mechanism for Local Differential Privacy

### v2

@@ Abstract.
With the increasing importance of data privacy, Local Differential Privacy (LDP) has recently become a strong measure of privacy for protecting each user’s privacy from data analysts without relying on a trusted third party. In this paper, we consider the problem of high-utility differentially private release. Given a domain of finite integers $\{1,2,\cdots,N\}$ and a distance-defined utility function, our goal is to design a differentially private mechanism that releases an item with the global expected error as small as possible. The most common LDP mechanism for this task is the Generalized Randomized Response (GRR) mechanism that treats all candidates equally except for the true item. In this paper, we introduce Bipartite Randomized Response mechanism (BRR), which adaptively divides all candidates into two parts by utility rankings given priori item. In the local search phase, we confirm how many high-utility candidates to be assigned with high release probability as the true item, which gives the locally optimal bipartite classification of all candidates. For preserving LDP, the global search phase uniformly selects the smallest number of dynamic high-utility candidates obtained locally. In particular, we give explicit formulas on the uniform number of dynamic high-utility candidates. The global expected error of our BRR is always no larger than the GRR, and can offer a decrease with a small and asymptotically exact factor. Extensive experiments demonstrate that BRR outperforms the state-of-the-art methods across the standard metrics and datasets.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/zpbh/BipartiteRR.
@@ 1. Introduction
(a) Quality loss $(QLoss)$
(b) $Q_{g}(BRR)$ / $Q_{g}(GRR)$
One foundational mechanism in LDP is the Randomized Response (RR) mechanism introduced by Warner (Warner, 1965). While RR is effective for protecting binary data, its applicability is limited to scenarios where the answer domain consists of only two values. To address this, the Generalized Randomized Response mechanism (GRR, also known as $k$-$R$) (Kairouz et al., 2016b) is proposed. GRR extends the RR mechanism to accommodate larger answer domains consisting of $N$ possible candidates. Under GRR, the data publisher selects and reports the true value with a fixed probability $p$, while assigning an equal probability to each of the remaining N1N-1 values with probability $\frac{1-p}{N-1}$. This extension retains the privacy-preserving properties of RR while broadening its applicability to use cases involving multi-valued categorical data, such as user preferences. It can be used directly as a competitive mechanism for a range of tasks, including frequency estimation (Fang et al., 2023; Wang et al., 2019), data collection  (Ye et al., 2019), decision tree  (Ma et al., 2023), and deep learning  (Mahawaga Arachchige et al., 2020). The GRR mechanism is order optimal in the low privacy regime (large $\epsilon$) for a series of utility functions (Kairouz et al., 2016b). However, GRR has some limitations on the scenarios as follows:
@@
Let $X$ denote the space of true values and $Y$ the space of released values, where $Y$ is identical to $\mathcal{X}=\{1,2,,\ldots,N\}$. When perturbing a specific $x_{i}$, BRR assigns higher publishing probabilities to the $m$ candidates most close to $x_{i}$ (with higher utility). By adjusting the distribution of publishing probabilities, BRR increases the likelihood of publishing data that are closer to the true item while decreasing the likelihood for data that are farther. This reallocation mechanism renders the publishing probabilities more rational.
@@ 3.3. Determining Globally Uniform $m$ for BRR
Step 2: Search globally optimal $m=\min_{k}m^{(k)}$, cf. Line 9.
@@ 3.4. Explicit Formulas for $m$ and Utility
During the local search phase by Alg. 1, one can mainly mention the sign of the derivative $\frac{\partial Q}{\partial s_{i}}$ for search the optimal local $m$. For the symbol $l\prec l^{\prime}$ in the column labeled by $\lambda_{i}$, their involved arrays are denoted by $\lambda]$ and $\lambda^{\prime}]$, respectively, and the inequality $\frac{\partial Q}{\partial s_{i}}\Big|_{[\lambda]}<\frac{\partial Q}{\partial s_{i}}\Big|_{[\lambda^{\prime}]}$ holds true based on the setting $s_{1}=\ldots=s_{i-1}=e^{\epsilon},\ s_{i+1}=\ldots=s_{N}=1$. This indicates that the array $\lambda^{\prime}]$ achieves the non-negative sign on the partial derivatives no later than array $\lambda]$ and their locally optimal number of high-weighted candidates $m^{\prime}\leq m$. This claim will be discussed in Appendix A.1 from (Zhang et al., 2025).
@@
0: Utility matrix $\{\lambda^{(k)}_{j}\}_{N\times N}$ with $0\leq\lambda_{1}^{(k)}\leq\lambda_{2}^{(k)}\leq\cdots\leq\lambda_{N}^{(k)}<N$, privacy budget $\epsilon$
@@
3:if $\epsilon\geq 1$, then $m=m_{1}$ and go to Line 7
4:if NN is even then $i=\left\lfloor\frac{N}{e^{\frac{\epsilon}{2}}+1}+1\right\rfloor$, otherwise, $i=\left\lfloor\frac{\sqrt{e^{\epsilon}(N^{2}-1)+1}-N}{e^{\epsilon}-1}+1\right\rfloor$
5:if $i$ is odd then $m_{2}=i$, otherwise, $m_{2}=i+1$
@@
3) BRR satisfies $\epsilon$-differential privacy.
@@
The proofs are presented in Appendix A.1 from (Zhang et al., 2025). ∎
Theorem 2.
respectively, and their gap is
Proof.
The proof is presented in Appendix A.2 from (Zhang et al., 2025). ∎
Theorem 3.
$\lim\limits_{N\to\infty}\frac{Q_{g}(BRR,N)}{Q_{g}(GRR,N)}=\frac{7e^{\frac{\epsilon}{2}}+1}{4e^{\frac{\epsilon}{2}}\left(e^{\frac{\epsilon}{2}}+1\right)}.$
Proof.
The proof is presented in Appendix A.3 from (Zhang et al., 2025). ∎
4. Exploring the scenarios of BRR
4.1. Adapting BRR in Continuous Domain
The Laplace mechanism satisfies $\epsilon$-DP by adding random noise to the output of query function $q$ on a given dataset. The distribution of random noises is the Laplace distribution with the scale $\Delta q/\epsilon$. Then we can transform the continuous noise obfuscation into a discretized perturbation process. Specifically, the perturbation interval defined by the scalar range of query function is equally divided and present a discrete domain consisting of $N$ discrete nodes,
$\mathcal{Y}=\{y_{1},y_{2},\dots,y_{n}\}.$
Then, the actual query value $x$ is firstly mapped to the nearest discrete value $y^{\ast}\in\mathcal{Y}$:
$y^{\ast}=\arg\mathop{\min}_{y_{i}\in\mathcal{Y}}|x-y_{i}|.$
Afterwards, the mapped value $y^{\ast}$ can be perturbed via the $BRR$ mechanism. Since the $BRR$ randomized response mechanism provides LDP guarantees, this method can flexibly improve the discretization accuracy by increasing the number of nodes while enforcing $\epsilon$-LDP.
The input of BRR already contains a certain level of noise during the mapping to discrete points. Specifically, since the original query value $x$ is mapped to the nearest point $y^{\ast}$ in the discrete set $\mathcal{Y}$, this mapping process introduces a deviation from the original $x$, causing data distortion. Therefore, to ensure the approximation accuracy, we can reduce the mapping error by increasing the number of discrete points $N$. As $N$ increases, the smaller intervals $L$ among the discrete points become finer, making the mapped value $y^{\ast}$ closer to the original data $x$, thus maintaining higher data accuracy.
For the vector case of Laplace mechanism, we can also develop the discrete domain in each dimension of the query range. More discussions will be demonstrated in Section 5.6.
4.2. BRR in General Discrete Domain
For the one-dimensional domain consisting of $N$ equidistant points, the points could be labeled by $1,2,\ldots,N$, in order. The linear transformation simply gives the BRR as above.
For stochastic distribution in the multi-dimensional discrete domain, or general metrics of data utility, one can perform the search in terms of Alg. 2 with necessary adjustments on the ranking of $\lambda_{i}$ and the sign condition of partial derivatives of utility.
Theorem 4.
(a)   $N=20$
(b)   $N=40$
(c)   $N=60$
(d)   $N=80$
(e)   $N=100$
5. Experimental Evaluations
In this section, the experimental evaluations consist of two parts:
  • Applications in continuous domains for machine learning (Sections 5.4 to 5.6). We evaluate the effect of applying BRR to machine learning for centralized DP, replacing Laplace mechanism in private GBDT (Li et al., 2020), SGD (Abadi et al., 2016), and vector perturbation for DNN, cf. (Wang et al., 2020).
5.1. Quality loss comparisons to GRR and Exponential mechanisms
As we know, the data utility is negatively related to the distance, like Euclidean and Jaccard distances. Following Section 3, we define Euclidean distance as the data utility (loss) $\lambda$ here. Then we conduct the experiments in the equidistant points domain $\mathcal{X}=\{1,2,\cdots,N\}$ with varying $N$. The standard metrics for comparisons among BRR, GRR and Exponential mechanisms include $Q_{g}$ ratios and $QLoss$, where
with the diameter normalization of domain $\mathcal{X}$.
5.2. BRR with Jaccard-Similarity Utility
Following the setting of Section 3, we conduct the experiments in the equidistant points domain $\mathcal{X}=\{1,2,\cdots,N\}$ with varying $N$. We denote data utility $\lambda$ by the pairwise generalized Jaccard similarity coefficient (Tinimoto coefficient), for any candidate $y$, given the true point $x$,
(13) $\lambda(x,y)=\frac{xy}{x^{2}+y^{2}-xy}.$
(a)   $N=20$
(b)   $N=40$
(c)   $N=60$
(d)   $N=80$
(e)   $N=100$
(a)   Gowalla, 1000 locations
(b)   FourSquare, 3000 locations
(a)   Gowalla
(b)   FourSquare
5.3. BRR in Location-based Services
(a)   Bias
(b)   abalone
(c)   sup
(a)   $\epsilon=0.1$
(b)   $\epsilon=0.3$
(c)   $\epsilon=0.5$
(d)   $\epsilon=0.7$
(e)   $\epsilon=0.9$
(f)   $\epsilon=1$
5.4. BRR in Private GBDT
Now, we turn to the applications of BRR in continuous domains for machine learning. The current experiments are conducted by deploying BRR mechanism as an alternative to the Laplace mechanism for perturbing each leaf value in DPBoost by Li et al. (Li et al., 2020). Originally, for training DPBoost, the sensitivity of the noise added after each geometric leaf clipping is $\Delta q=\min(\frac{g_{l}^{\ast}}{1+\lambda},2g_{l}^{\ast}(1-\eta)^{t-1})$. In the training of new DP-GBDT called DPBoost-BRR due to the above alternative, the perturbation range of leaf clipping should be [c,c][-c,c] with $c=\Delta q/2$. Then the range [c,c][-c,c] is equally divided for suitable $L$ that gives $n=\frac{2c}{L}+1$ discrete equidistant points. The leaf node values are first adjusted to the nearest candidate. Finally, applying the BRR mechanism gives the perturbed leaf values.
5.5. BRR in Private SGD
(a)   US Census
(b)   Bank Marketing
5.6. BRR for Vector Perturbation in DNN
6. Related Work
6.1. Advancements in LDP Mechanisms
6.2. Applications of LDP
LDP in Location-based Services. Wang et al.  (Wang et al., 2022) proposed a staircase randomized response mechanism named L-SRR that introduced the hierarchical coding scheme. Zhang et al. (Zhang et al., 2024) developed a regionalized LDP mechanism called DPIVE with personalized utility sensitivities for LBSs by incorporating expected inference error and geo-indistinguishability. Wu et al.  (Wu et al., 2020) investigated noise injection to protect privacy in distributed machine learning in multi-data-owner environments. Gutiérrez et al.  (Gutiérrez et al., 2024) introduced a privacy-preserving framework for fog computing environments while maintaining model accuracy. Cao et al.  (Cao et al., 2021) investigated data poisoning attacks on LDP protocols and proposed defense methods.
7. Conclusion
This paper proposes a novel adaptive LDP mechanism, the Bipartite Randomized Response (BRR), which aims to balance the trade-off between privacy and data utility. In particular, the GRR mechanism treats all non-true items equally, which weakens the data utility in some scenarios. To address this limitation, BRR optimizes the allocation of release probabilities by local and global search phases. This optimization enhances data utility while preserving privacy level. We provide a series of theoretical results for the optimal number $m$ of high-weighted items and asymptotically exact ratio of BRR to GRR or global expected errors in the equidistant points domains. various empirical valuations verify that our BRR approach outperforms the state-of-the-art methods.
For the future work, we will investigate some reformed Bipartite Random Response mechanisms for minimizing the global expected error in general discrete domains under uniform and prior distributions.
Appendix A Appendix
A.1. Proof of in Theorem 1
The proof begins with the illustration by Fig. 3. For each adjacent priori number $k$, the distances for all candidates are sorted in an ascending order. For any two adjacent rows with priori numbers $k$ and $k^{\prime}=k-1$, respectively, we compare their sorted distance arrays (rows), $\lambda]$ and $\lambda^{\prime}]$. Then there exists a single $j_{0}$, such that $\lambda_{1}^{\prime}=\lambda_{1},\ \lambda_{2}^{\prime}=\lambda_{2},\ \cdots,\ \lambda_{j_{0}}^{\prime}=\lambda_{j_{0}}$, while for the remaining $N-j_{0}$ distances, $\lambda_{j_{0}+1}^{\prime}=\lambda_{j_{0}+1}+1,\ \lambda_{j_{0}+2}^{\prime}=\lambda_{j_{0}+2}+1,\ \cdots,\ \lambda_{N}^{\prime}=\lambda_{N}+1$. Such a number $j_{0}$ is indicated by the column number involving a gray ellipse area between the two rows.
Now we consider to increase the probability weight $s_{i}$ in order. We first consider the case $i\leq j_{0}$. When considering the weight increase on $s_{i}$, we have fixed $s_{1}=\cdots=s_{i-1}=e^{\epsilon}$ for both rows and
$\begin{split}\sum_{j}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}=&\sum_{j=1}^{j_{0}}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}\\ =&\sum_{j=1}^{j_{0}}(\lambda_{i}-\lambda_{j})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}-(\lambda_{j}+1))s_{j}\\ =&\sum_{j=1}^{j_{0}}(\lambda_{i}-\lambda_{j})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}-\lambda_{j})s_{j}-\sum_{j=j_{0}+1}^{N}s_{j}\\ =&\sum_{j}(\lambda_{i}-\lambda_{j})s_{j}-\sum_{j_{0}+1}^{N}s_{j}\\ \leq&\sum_{j}(\lambda_{i}-\lambda_{j})s_{j},\end{split}$
@@ A.1. Proof of in Theorem 1
$\begin{split}\sum_{j}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}=&\sum_{j=1}^{j_{0}}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}\\ =&\sum_{j=1}^{j_{0}}(\lambda_{i}+1-\lambda_{j})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}+1-(\lambda_{j}+1))s_{j}\\ =&\sum_{j=1}^{j_{0}}(\lambda_{i}-\lambda_{j})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}-\lambda_{j})s_{j}+\sum_{j=1}^{j_{0}}s_{j}\\ =&\sum_{j}(\lambda_{i}-\lambda_{j})s_{j}+\sum_{j=1}^{j_{0}}s_{j}\\ >&\sum_{j}(\lambda_{i}-\lambda_{j})s_{j}.\end{split}$
@@
Based on Fig. 3, if the threshold value $i$ for changing the sign of the partial derivative appears before (no more than) $j_{0}$ (in the red region), the lower row will always reach the threshold before the upper one for any two adjacent rows. In this case, starting from the midpoint $N/2$ as the prior, the bottom-most row achieves the smallest $m$. On the other hand, if the threshold appears after (larger than) $j_{0}$ (in the green region), the upper row $k$ will reach the threshold before the lower $k^{\prime}$. In this scenario, starting from the edge points $1$ or $N$ as the prior, the top-most row achieves the smallest $m$. Fig. 3 depicts the situation where $N$ is even; the case for odd $N$ is similar and will not be elaborated further.
@@
(14) $\begin{split}&\sum_{j=1}^{i-1}(\lambda_{i}-\lambda_{j})e^{\epsilon}+\sum_{j=i+1}^{N}(\lambda_{i}-\lambda_{j})1\\ =&\frac{(1+i-1)(i-1)}{2}e^{\epsilon}-\frac{(1+N-i)(N-i)}{2}\\ =&\frac{i(i-1)}{2}e^{\epsilon}-\frac{(N-i)^{2}}{2}-\frac{N-i}{2}\\ =&\frac{1}{2}(e^{\epsilon}-1)i^{2}+\left(N-\frac{1}{2}e^{\epsilon}+\frac{1}{2}\right)i-\frac{1}{2}N^{2}-\frac{1}{2}N,\end{split}$
@@
(16) $\begin{split}&\sum_{j=1}^{i-1}(\lambda_{i}-\lambda_{j})e^{\epsilon}+\sum_{j=i+1}^{N}(\lambda_{i}-\lambda_{j})\times 1\\ =&\frac{i-1}{2}e^{\epsilon}+(1+2+\cdots+\frac{i-1}{2}-1)e^{\epsilon}\times 2\\ &-(1+2+\cdots+\frac{N-1}{2}-\frac{i-1}{2})\times 2\\ =&\frac{(e^{\epsilon}-1)}{4}i^{2}+\frac{N-(e^{\epsilon}-1)}{2}i+\frac{(e^{\epsilon}-1)-2N-N^{2}+1}{4},\end{split}$

### v3

@@ Abstract.
With the increasing importance of data privacy, Local Differential Privacy (LDP) has recently become a strong measure of privacy for protecting each user’s privacy from data analysts without relying on a trusted third party. In this paper, we consider the problem of high-utility differentially private release. Given a domain of finite integers $\{1,2,\cdots,N\}$ and a distance-defined utility function, our goal is to design a differentially private mechanism that releases an item with the global expected error as small as possible. The most common LDP mechanism for this task is the Generalized Randomized Response (GRR) mechanism that treats all candidate items equally except for the true item. In this paper, we introduce Bipartite Randomized Response mechanism (BRR), which adaptively divides all candidate items into two parts by utility rankings given priori item. In the local search phase, we confirm how many high-utility candidates to be assigned with high release probability as the true item, which gives the locally optimal bipartite classification of all candidates. For preserving LDP, the global search phase uniformly selects the smallest number of dynamic high-utility candidates obtained locally. In particular, we give explicit formulas on the uniform number of dynamic high-utility candidates. The global expected error of our BRR is always no larger than the GRR, and can offer a decrease with a small and asymptotically exact factor. Extensive experiments demonstrate that BRR outperforms the state-of-the-art methods across the standard metrics and datasets.
Local Differential Privacy, Bipartite Randomized Response, Utility
copyright: acmlicensedjournalyear: 2026doi: $X$.XXXXconference: ACM Conf.; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Privacy-preserving protocolsccs: Security and privacy Data anonymization and sanitization
@@ 1. Introduction
(a) GRR mechanism (Kairouz et al., 2016b)
(b) BRR mechanism
Figure 1. Probability mass function (PMF) for $GRR$ and $BRR$.
One foundational mechanism in LDP is the Randomized Response (RR) mechanism introduced by Warner (Warner, 1965). While RR is effective for protecting binary data, its applicability is limited to scenarios where the answer domain consists of only two values. To address this, the Generalized Randomized Response mechanism (GRR, also known as $k$-$R$) (Kairouz et al., 2016b) is proposed. GRR extends the RR mechanism to accommodate larger answer domains consisting of $N$ possible candidates as shown in Fig. 1(a). Under GRR, the data publisher selects and reports the true value with a fixed probability $p$, while assigning an equal probability to each of the remaining N1N-1 values with probability $\frac{1-p}{N-1}$. This extension retains the privacy-preserving properties of RR while broadening its applicability to use cases involving multi-valued categorical data, such as user preferences. It can be used directly as a competitive mechanism for a range of tasks, including frequency estimation (Fang et al., 2023; Wang et al., 2019), data collection  (Ye et al., 2019), decision tree  (Ma et al., 2023), and deep learning  (Mahawaga Arachchige et al., 2020). The GRR mechanism is order optimal in the low privacy regime (large $\epsilon$) for a series of utility functions (Kairouz et al., 2016b). However, GRR has some limitations on the scenarios as follows:
@@
Let $X$ denote the space of true values and $Y$ the space of released values, where $Y$ is identical to $\mathcal{X}=\{1,2,,\ldots,N\}$. When perturbing a specific $x_{i}$, BRR assigns higher publishing probabilities to the $m$ candidates most close to $x_{i}$ (with higher utility) as shown in Fig. 1(b). By adjusting the distribution of publishing probabilities, BRR increases the likelihood of publishing data that are closer to the true item while decreasing the likelihood for data that are farther. This reallocation mechanism renders the publishing probabilities more rational.
@@ 3.3. Determining Globally Uniform $m$ for BRR
Step 2: Search globally uniform $m=\min_{k}m^{(k)}$, cf. Line 9.
@@ 3.4. Explicit Formulas for $m$ and Utility
During the local search phase by Alg. 1, one can mainly mention the sign of the derivative $\frac{\partial Q}{\partial s_{i}}$ for search the optimal local $m$. For the symbol $l\prec l^{\prime}$ in the column labeled by $\lambda_{i}$, their involved arrays are denoted by $\lambda]$ and $\lambda^{\prime}]$, respectively, and the inequality $\frac{\partial Q}{\partial s_{i}}\Big|_{[\lambda]}<\frac{\partial Q}{\partial s_{i}}\Big|_{[\lambda^{\prime}]}$ holds true based on the setting $s_{1}=\ldots=s_{i-1}=e^{\epsilon},\ s_{i+1}=\ldots=s_{N}=1$. This indicates that the array $\lambda^{\prime}]$ achieves the non-negative sign on the partial derivatives no later than array $\lambda]$ and their locally optimal number of high-weighted candidates $m^{\prime}\leq m$. This claim will be discussed in the proof of Theorem 1.
@@
0: Privacy budget $\epsilon$, equidistant points count NN
@@
2:if $\epsilon\geq 1$, then $m=m_{1}$ and go to Line 6
3:if NN is even then $i=\left\lfloor\frac{N}{e^{\frac{\epsilon}{2}}+1}+1\right\rfloor$, otherwise, $i=\left\lfloor\frac{\sqrt{e^{\epsilon}(N^{2}-1)+1}-N}{e^{\epsilon}-1}+1\right\rfloor$
4:if $i$ is odd then $m_{2}=i$, otherwise, $m_{2}=i+1$
@@
3) BRR satisfies $\epsilon$-local differential privacy.
@@
We begin with the illustration by Fig. 3. For each adjacent priori number $k$, the distances for all candidates are sorted in an ascending order. For any two adjacent rows with priori numbers $k$ and $k^{\prime}=k-1$, respectively, we compare their sorted distance arrays (rows), $\lambda]$ and $\lambda^{\prime}]$. Then there exists a single $j_{0}$, such that $\lambda_{1}^{\prime}=\lambda_{1},\ \lambda_{2}^{\prime}=\lambda_{2},\ \cdots,\ \lambda_{j_{0}}^{\prime}=\lambda_{j_{0}}$, while for the remaining $N-j_{0}$ items, $\lambda_{j_{0}+1}^{\prime}=\lambda_{j_{0}+1}+1,\ \lambda_{j_{0}+2}^{\prime}=\lambda_{j_{0}+2}+1,\ \cdots,\ \lambda_{N}^{\prime}=\lambda_{N}+1$. Such a number $j_{0}$ is indicated by the column number involving a gray ellipse area between the two rows.
Now we consider to increase the probability weight $s_{i}$ in order. We first consider the case $i\leq j_{0}$. When considering the weight increase on $s_{i}$, we have fixed $s_{1}=\cdots=s_{i-1}=e^{\epsilon}$ for both rows, and for the numerator part of (9) about $\frac{\partial Q}{\partial s_{i}}$,
$\begin{split}\sum_{j}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}=&\sum_{j=1}^{j_{0}}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}\\ =&\sum_{j=1}^{j_{0}}(\lambda_{i}-\lambda_{j})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}-(\lambda_{j}+1))s_{j}\\ =&\sum_{j}(\lambda_{i}-\lambda_{j})s_{j}-\sum_{j_{0}+1}^{N}s_{j}\leq\sum_{j}(\lambda_{i}-\lambda_{j})s_{j},\end{split}$
@@
$\begin{split}\sum_{j}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}=&\sum_{j=1}^{j_{0}}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}^{\prime}-\lambda_{j}^{\prime})s_{j}\\ =&\sum_{j=1}^{j_{0}}(\lambda_{i}+1-\lambda_{j})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}+1-(\lambda_{j}+1))s_{j}\\ =&\sum_{j=1}^{j_{0}}(\lambda_{i}-\lambda_{j})s_{j}+\sum_{j=j_{0}+1}^{N}(\lambda_{i}-\lambda_{j})s_{j}+\sum_{j=1}^{j_{0}}s_{j}\\ =&\sum_{j}(\lambda_{i}-\lambda_{j})s_{j}+\sum_{j=1}^{j_{0}}s_{j}>\sum_{j}(\lambda_{i}-\lambda_{j})s_{j}.\end{split}$
@@
Based on Fig. 3, if the threshold value $i$ for changing the sign of the partial derivative appears before (no more than) $j_{0}$ (in the red region), the lower row will always reach the threshold before the upper one for any two adjacent rows. In this case, starting from the midpoint $N/2$ as the prior, the bottom-most row achieves the smallest $m$. On the other hand, if the threshold appears after (larger than) $j_{0}$ (in the green region), the upper row $k$ will reach the threshold before the lower $k^{\prime}$. In this scenario, starting from the extreme points $1$ or $N$ as the prior, the top-most row achieves the smallest $m$. Fig. 3 depicts the situation where $N$ is even; the case for odd $N$ is similar and will not be elaborated further.
@@
(12) $\begin{split}&\sum_{j=1}^{i-1}(\lambda_{i}-\lambda_{j})e^{\epsilon}+\sum_{j=i+1}^{N}(\lambda_{i}-\lambda_{j})1\\ =&\frac{(1+i-1)(i-1)}{2}e^{\epsilon}-\frac{(1+N-i)(N-i)}{2}\\ =&\frac{1}{2}(e^{\epsilon}-1)i^{2}+\left(N-\frac{1}{2}e^{\epsilon}+\frac{1}{2}\right)i-\frac{1}{2}N^{2}-\frac{1}{2}N,\end{split}$
@@
(14) $\begin{split}&\sum_{j=1}^{i-1}(\lambda_{i}-\lambda_{j})e^{\epsilon}+\sum_{j=i+1}^{N}(\lambda_{i}-\lambda_{j})\times 1\\ =&\frac{(e^{\epsilon}-1)}{4}i^{2}+\frac{N-(e^{\epsilon}-1)}{2}i+\frac{(e^{\epsilon}-1)-2N-N^{2}+1}{4},\end{split}$